OpenStudy (anonymous):

Write an equation of a line with a slopem=1/2 and including points (4,-1)

OpenStudy (anonymous):

-2x+3 2x+3 y=1/2x+3 y=1/2x-3

OpenStudy (hba):

Use the formula, \[y-y_1=m(x-x_1 )\]

OpenStudy (anonymous):

ok I'll try and see what I get thansk!

OpenStudy (anonymous):

How owuld I set that up though?

OpenStudy (hba):

(x1,y1)= (4,-1) m=1/2 Now solve :)

OpenStudy (anonymous):

Okay this is what I haveY-(-1)=1/2(x-4)?